OpenStudy (samsan9):

find the center, vertices, foci, eccentricity, length of major axis, and length of minor axis for (x-5)^2/4+(y+3)^2/16=1

OpenStudy (mathmate):

Do you know what conic it is?

OpenStudy (samsan9):

sort of its like hyperbolas, parabolas, ellipses, and i believe circles

OpenStudy (mathmate):

Can you tell which one from your text book?

OpenStudy (samsan9):

it is all of them so far and i have to draw the graph as well :/

OpenStudy (mathmate):

Each conic has a characteristic and a standard form. If you study conics, the first thing you need to know is to tell which one it is! For example, Equation of every circle can be reduced to: \(\large (x-a)^2+(y-b)^2=r^2\)

OpenStudy (mathmate):

Does it look like a circle?

OpenStudy (samsan9):

I think its an ellipse since someone told me that this is what they usually look like if i remember correctly

OpenStudy (mathmate):

Yes, it is an ellipse, with the standard equation:

OpenStudy (samsan9):

\[\frac{ (x-5)^2 }{ 4 }+\frac{ (y+3)^2 }{ 16 }=1\]

OpenStudy (mathmate):

\(\large \frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=-1\)

OpenStudy (mathmate):

Oops: \(\large \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\)

OpenStudy (mathmate):

where (h,k) is the centre of the ellipse, a and be are the axes. If a>b, the major (longer) axis is horizontal, and if a<b, the major axis is vertical.

OpenStudy (mathmate):

ok so far?

OpenStudy (mathmate):

a and b are the distances from the centre to the vertices.

OpenStudy (mathmate):



OpenStudy (mathmate):

Can you find h,k, a,b from the given equation, in comparison with the general equation?

OpenStudy (samsan9):

h=5 and k=-3 i think and the a=2 and b=4

OpenStudy (samsan9):

am i right?

OpenStudy (mathmate):

That is correct!

OpenStudy (mathmate):

So a is x-distance between centre and vertex, b is y-distance of same.

OpenStudy (mathmate):

Do you know the definition of major and minor axes?

OpenStudy (mathmate):

*definitions

OpenStudy (samsan9):

major is the square root of the lower number times 2 and major is the squareroot of the highest number times 2?

OpenStudy (mathmate):

exactly, in other words, major axis is 2a or 2b, whichever is greater. Similarly for minor.

OpenStudy (mathmate):

How about eccentricity?

OpenStudy (samsan9):

that i can't remember :/

OpenStudy (mathmate):

eccentricity = sqrt(1-(minor axis/major axis)^2)

OpenStudy (samsan9):

\[\sqrt{1-4/8}?\]

OpenStudy (mathmate):

(4/8)squared!

OpenStudy (samsan9):

so -1-1/2? i am sorry i dont get this part

OpenStudy (mathmate):

http://en.wikipedia.org/wiki/Eccentricity_(mathematics) in case you need a reference. eccentricity = sqrt(1-(4/8)^2)=sqrt(1-1/4)=sqrt(3/4)=sqrt(3)/2

OpenStudy (mathmate):

sorry, afk.

OpenStudy (samsan9):

oh ok

OpenStudy (samsan9):

what about the vertices and foci?

OpenStudy (mathmate):

Vertices: (h \(\pm a, k \pm b)\) c=sqrt(major^2-minor^2) foci: \((h \pm c, k)\) for a>b \((h, k \pm c)\) for a<b